Consider the following accounting statement for a small economy:

Compensation of employees $10,000 Consumption of fixed capital $100 Corporate income taxes $200 Dividends $350 Government purchases $1,000 Gross private domestic investment $1,000 Indirect business taxes $400 Interest $600 Net exports -$1,000 Net foreign factor income earned in the United States $50 Personal consumption expenditures $12,000 Personal taxes $1,050 Proprietors' income $800 Rents $200 Social Security contributions $500 Transfer payments $200 Undistributed corporate profits $300.

In this economy, GDP is equal to what amount?
Answer #1

ANswer

GDP=C+I+G+NX

C=Personal consumption expenditures $12,000

I=Gross private domestic investment $1,000

G= Government purchases $1,000

NX=Net exports -$1,000

GDP=12000+1000+1000+(-1000)=$13000

the GDP is $13000

==========

  
GDP=Compensation of employees $10,000+Consumption of fixed capital $100+Corporate income taxes $200 +Dividends $350 +Indirect business taxes $400 +Interest $600+Net foreign factor income earned in the United States $50+Proprietors' income $800+ Rents $200+Undistributed corporate profits $300
=10000+100+200+350+400+600+50+800+200+300
=13000